Circular No. 7911

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 K4 (NEAT) Additional observations (cf. IAUC 7909) and the following preliminary

parabolic orbital elements appear on MPEC 2002-K69:

T = 2002 Aug. 8.013 TT  ω = 32.^{o}810
Ω = 308.814 }2000.0
q = 2.74229 AU  i = 95.939

SUPERNOVA 2002cv IN NGC 3190 P. Meikle and S. Mattila, Imperial College, London, report that in-

frared magnitudes of SN 2002cv obtained by A. Glasse at the U.K. Infrared Telescope (UKIRT) on May 17.3 UT (H = 14.3, K = 13.9) indicate a brightening since May 13.8 (cf. IAUC 7901). Infrared spectra (range 0.83–1.35 and 1.98–2.5 μm) were obtained by J. Buckle at the UKIRT on May 22.3 and 23.3. A spectrum covering bands I and J (and of fairly low S/N) shows a continuum declining by ∼ 30 percent from the short- to long-wavelength limits. No strong features are present. The K-band spectrum (with good S/N) shows a number of broad, prominent emission/absorption features superimposed on a continuum that declines toward longer wavelengths. The presence of these features means that the infrared spectrum is not that of a type-II supernova. An infrared spectrum (range 0.83–2.5 μm) of the type-Ia SN 2002bo (IAUC 7846), also in NGC 3190, was obtained at the UKIRT by A. Adamson on Apr. 3.3, ∼ 10 days after maximum light (B-band); by reddening the SN 2002bo infrared spectrum relative to that of SN 2002cv, by an absolute V -band extinction of A(V) ∼ 6 (using the extinction law of Cardelli et al. 1989, Ap.J. 345, 245), the two infrared spectra become quite similar in appearance, with SN 2002cv being ∼ 0.4 mag brighter than the dereddened SN 2002bo. Meikle and Mattila add: “We conclude that the absolute (d = 22 Mpc) dereddened magnitudes of SN 2002cv (cf. Meikle 2000, MNRAS 314, 782) and the appearance of its infrared spectra indicate a type-Ia event with maximum light (B-band) around May 23. Given the earlier, higher-temperature epoch of the SN 2002cv infrared spectra, compared with SN 2002bo, it is likely that the value A(V) toward SN 2002cv actually exceeds 6, making this probably the highest ever recorded for an identified type-Ia supernova. Unlike SN 2002bo, the SN 2002cv I-band spectrum does not show strong Ca II triplet emission/absorption. However, weak Ca II has been observed in some type-Ia supernovae such as the peculiar SN 1991T at maximum light.”

2002 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green